# Verification and Refinement of Natural Language Explanations through LLM-Symbolic Theorem Proving

### v2

@@ 4 Empirical Evaluation
To assess the effectiveness of integrating LLMs with external feedback for refinement purposes, we evaluated the Explanation-Refiner on two NLI tasks: textual entailment and multiple-choice question answering. For textual entailment, given a premise $p_{i}$ and an explanatory sentence as explanation $E_{i}$, the goal is to determine whether $p_{i}$ and $E_{i}$ together entail a hypothesis $h_{i}$. In the multiple-choice question answering task, there is a question $q$ accompanied by a set of candidate answers $C=\{c_{1},c_{2},...,c_{n}\}$, with $c_{i}$ identified as the correct answer. We convert $q$ and the correct answer $c_{i}$ as the hypothesis $h_{i}$. Explanatory facts serve as evidence supporting $c_{i}$ as the correct answer and are denoted as $E_{i}$, with the question’s context sentence as the premise $p_{i}$.
@@ 4.1 Datasets
We adopted three different NLI datasets for evaluation: e-SNLI (Camburu et al., 2018), QASC (Khot et al., 2019), and WorldTree (Jansen et al., 2018), using a total of 300 samples selected via the sampling strategy defined in (Valentino et al., 2021), which maximises representativeness and mutual exclusivity across syntactic and semantic features expressed in the datasets. e-SNLI is a crowd-sourced dataset typically used as a benchmark for textual entailment. It comprises one premise sentence, one explanatory sentence, and one hypothesis sentence per sample. QASC and WorldTree, on the other hand, are datasets designed for multiple-choice question answering within the scientific domain. QASC includes two explanatory sentences for each correct answer. WorldTree is the most complex of the three, featuring between 1 and 16 explanatory sentences in each sample.
4.2 Theorem Prover
We adopted Isabelle/HOL (Nipkow et al., 2002) proof assistant as the theorem prover. To integrate this theorem prover as a real-time verification tool with LLMs, we employ a Python client (Shminke, 2022) as TCP (Transmission Control Protocol) client to configure Isabelle/HOL as a server. This enables the communication of the constructed theory files and the extraction of the response messages from Isabelle.
@@ 4.3 Models
We conducted experiments using five LLMs within the proposed framework. The models include two open-sourced models: Llama2-70b (Touvron et al., 2023) and Mixtral-8x7b (Jiang et al., 2024a), as well as Mistral-Small (mistral-small-latest) (Mistral AI, 2024), GPT-Turbo-3.5 (gpt-3.5-turbo) (Brown et al., 2020), and GPT-4 (gpt-4-0613) (OpenAI, 2023). Temperature settings were adjusted to 0 for GPT-Turbo-3.5 and GPT-4, and to 0.01 for Llama2-70b, Mixtral-8x7b, and Mistral-Small, aiming to achieve both determinism in the output and effective code generation for theorem prover.
@@ 4.4 Results
Providing feedback on the exact failure step from an external theorem prover effectively guides LLMs in continuously verifying and refining explanations in NLI tasks.
@@
Explanation length/complexity impacts formalisation.
The e-SNLI dataset, which includes only a single explanatory sentence per example, shows the best overall performance. In contrast, the multiple-choice question answering datasets, QASC and WorldTree, exhibit comparatively lower performance. QASC typically contains 2 explanatory sentences, while WorldTree ranges from 1 to 16 sentences. As the number of explanatory sentences increases, so does the complexity of the logical reasoning required. The WorldTree dataset, in particular, poses the greatest challenge due to its demand for multi-hop inference strategies. Models show lower refining performance in WorldTree when compared to e-SNLI and QASC, with only 3%, 5%, and 6% of Llama-70b, Mixtral-8x7b, and Mistral-Small explanations being refined in WorldTree. Meanwhile, 29% and 37% of explanations are refined by GPT-3.5-Turbo and GPT-4 in WorldTree, respectively. This process involves synthesising multiple explanatory sentences to fulfill sub-goals, which must then be integrated to meet the overall hypothesis goal.
Iterative and categorical refinement can monotonically reduce syntax errors in responses generated by LLMs.
@@ 4.5 Ablation Study
We conducted an ablation study to further evaluate and disentangle the impact of autoformalisation on performance. To this end, we adopted GPT-4 exclusively for the autoformalisation component, while retaining the original models for explanation refinement and proof strategy generation. As shown in Figure 7, integrating GPT-4 for autoformalisation led to a significant increase in the number of explanations successfully refined across all models. For instance, Llama2-70b with GPT-4 as the formalisation component refined explanations from 7% to 65% in the e-SNLI dataset. For the multiple-choice question answering dataset, GPT-3.5-Turbo showed a relatively smaller increase from 44% to 48% and from 29% to 34%. Despite these improvements, a performance gap persists between GPT-4 and the other models, which is attributed to GPT-4’s superior symbolic reasoning capabilities required for explanation refinement from the identified logical errors.
Explanations are progressively made more complete and consistent through iterative symbolic refinement.
In order to deliver step-wise logical consistency, explanations need to be made complete and self-contained, leading to the introduction of additional explanatory sentences, leading to an increase in the total number of suggested proof steps. Therefore, we further evaluated how the proof steps vary when the total number of suggested proof steps increases contrasting both refined and unrefined cases. Figure 8 illustrates this trend. In general, all models show a positive trend, as the total suggested proof steps increase, the average number of proof steps processed by the proof assistant also increases. Models like Mistral-Small and GPT-3.5-Turbo tend to suggest more proof steps to accomplish the logical goal, which can result in some redundant steps, such as the significant pulse shown in Figure 8(c). For unrefined explanations, as shown in Figure 8(d), 8(e) and 8(f), the progression is steadier but retains a positive trend, where the models generally suggest more proof steps in response to the additional explanatory sentences introduced to correct a logical error identified from the erroneous step. We also conducted experiments on the relationship of average number of successfully processed explanatory sentences in one proof against total planned explanatory sentences in a suggest proof in appendix A.3. Examples of refined and unrefined explanations can be found in Appendix A.5.
Figure 9: Human evaluation of refined explanations in terms of factuality and triviality. Factually Correct indicates the percentage of explanation sentences that are correct in terms of commonsense knowledge. Not Trivial states the percentage of explanations that contain only explanatory sentences that repeat or paraphrase the premise and/or the hypothesis.
@@ 4.6 Factual Errors and Trivial Explanations
In addition to evaluating the logical validity of explanations, we also conducted a human evaluation of the refined explanations, considering factual correctness and explanation triviality, for the two best-performing models (GPT-3.5-Turbo and GPT-4). This evaluation focused on two questions: “Are the refined explanatory sentences factually correct?” and “Is the explanation trivial, merely repeating or paraphrasing the content of the premise and hypothesis to achieve logical validity?”. As illustrated in Figure 9, our findings indicate that all refined explanations in the e-SNLI and WorldTree datasets are consistent with commonsense knowledge. In the QASC dataset, 2.27% and 1.82% of the explanation refined by GPT-3.5-Turbo and GPT-4 contain sentences misaligned with true world knowledge. We found that the majority of these errors result from over-generalisation, such as the sentence All tetrapods are defined to have four limbs, which inaccurately includes snakes.
@@ 5.1 LLMs Self-Refinement from External Feedback
Self-refinement of LLMs has demonstrated promising effectiveness in generating faithful and trustworthy responses (Pan et al., 2023b). The use of external feedback to guide LLMs has been extensively studied (Olausson et al., 2024a; Yu et al., 2023; Akyurek et al., 2023). Previous work such as Peng et al. (2023) and Li et al. (2024) have employed facts retrieved from external knowledge bases as sources of feedback, while Paul et al. (2024) developed a critic model to provide feedback for reasoning refinement. Additionally, Nathani et al. (2023) have explored the use of feedback models for automated feedback generation. Various works have also investigated tasks related to code generation (Chen et al., 2023; Olausson et al., 2024b) and the creation of either synthetic or expert-written logical natural language expressions (Olausson et al., 2023). Quan et al. (2024) leverage a differentiable logic reasoner to verify and refine explanations through abductive reasoning , enhancing the logical consistency of explanations in ethical NLI tasks based on the solver’s output. This paper focuses on the automated refinement of natural language sentences created by human annotators. It verifies and refines the logical validity of these human-generated sentences using detailed external feedback, which can identify the exact erroneous steps to effectively refine logical errors in the explanatory sentences.
@@ 5.2 Autoformalisation
Autoformalisation refers to the process of translating natural language descriptions into symbolic representations. Research in this area has included the formalisation of mathematical proofs (Cunningham et al., 2022; Wu et al., 2022; First et al., 2023), and efforts to transform natural language sentences into logical forms using LLMs (Pan et al., 2023a; Olausson et al., 2023; Jiang et al., 2024b; Dalal et al., 2024). However, contextual information is frequently lost when sentences are translated in these logical frameworks. To mitigate semantic loss during the transformation process, we leverage Neo-Davidsonian event semantics, which aims to maximise the preservation sentence-level content. This representation paradigm can facilitate a more systematic content-preserving translation to logical forms, which is more independent from particular choices of representation schemas.
5.3 Conclusion
In this work, we present a novel neuro-symbolic framework that automatically verifies and refines natural language explanations using iterative refinement cycles between LLMs and theorem provers (TPs). We have conducted extensive experiments on both textual entailment and multiple-choice question answering tasks, demonstrating that the proposed method, Explanation-Refiner, effectively enhances the logical validity of such human-annotated explanations. We investigated the model’s performance from simple to complex explanatory/sentence structures and introduced a method to prevent the loss of semantic information in autoformalisation tasks with error correction. Furthermore, we investigated the iterative model at each iteration to evaluate how the explanation is refined at each cycle. In future work, we aspire to enhance the framework’s robustness towards complex and unstructured explanations with fewer iterations required to improve the model’s efficiency.
5.4 Limitations
While this work have demonstrated significant improvements in terms of enhancing the logical consistency of explanations, the connection between improved logical consistency and AI safety still needs further investigation. While the concept of using formal solvers in conjunction with LLMs delivers a promise avenue to improve the consistency of reasoning within LLMs, these methodologies needs to be further developed and critically assessed as a mechanism which can provide guarantees of correctness, consistency and completeness within critical application domains.
@@ A.5 Examples of Explanation Refinement
A.5.1 e-SNLI Refined Examples
Table 1 shows an example from the e-SNLI dataset of how the explanation changes after each iteration. Figures 22, 23, and 24 illustrate the Isabelle/HOL theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
Table 2 along with Figures 25 and 26, and Table 3 with Figures 27, 28, and 29 are two more examples.
A.5.2 QASC Refined Examples
Table 4 shows an example from the QASC dataset of how the explanation changes after each iteration. Figures 30, 31 illustrate the Isabelle/HOL theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
Table 5 along with Figures 32 and 33, and Table 6 with Figures 34, 35 are two more examples.
A.5.3 WorldTree Refined Examples
Table 7 shows an example from the WorldTree dataset of how the explanation changes after each iteration. Figures 36, 37, 38, 39, 40, 41, 42 and 43 illustrate the Isabelle/HOL theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
Table 8 and Figures 44, 46, 47 and 48, as well as Table 9 with Figures 49, 50, 51, and 52, provide two more examples.
A.5.4 Unrefined Example
Table 10 and Table 11 shows an example from the WorldTree dataset that does not refine within 10 iterations and is not caused by a syntax error. The figures that follow show the detailed Isabelle/HOL theory code of the related iterations.
@@ 44      $iterations\leftarrow iterations+1$
e-SNLI Premise: A couple being romantic under the sunset.
Hypothesis: People are enjoying sundown.
A couple consists of people. Sunset is sundown. 0 Invalid
e-SNLI Premise: A couple being romantic under the sunset.
Hypothesis: People are enjoying sundown.
A couple consists of people. Sunset is sundown. Being romantic under the sunset implies enjoying the sundown. 1 Valid
Figure 25: The Isabelle theory code for table 2 iteration 0
Figure 26: The Isabelle theory code for table 2 iteration 1
Dataset Sentences Explanation Iteration Validity

### v4

@@ 4.1 Datasets
We adopted three different NLI datasets for evaluation: e-SNLI, QASC, and WorldTree, using a total of 300 samples selected via the sampling strategy defined in Valentino et al. (2021), which maximises representativeness and mutual exclusivity across syntactic and semantic features expressed in the datasets. For multiple-choice question answering, the task includes a question $q$ accompanied by a set of candidate answers $C=\{c_{1},c_{2},...,c_{n}\}$, with $c_{i}$ identified as the correct answer. To cast this problem into NLI, we simply convert $q$ and the correct answer $c_{i}$ into a hypothesis $h_{i}$. On the other hand, the question’s context, if present, is used to build the premise $p_{i}$.
@@ 4.2 Models
To integrate Isabelle/HOL as a real-time verification tool with LLMs, we employ a Python client (Shminke, 2022) which communicates with Isabelle/HOL as a server backend. This enables the communication of the constructed theory files and the extraction of the response messages from Isabelle. We conducted experiments using five LLMs within the proposed framework. The models include two open-sourced models: Llama2-70b (Touvron et al., 2023) and Mixtral-8x7b (Jiang et al., 2024a), as well as Mistral-small (mistral-small-latest) (Mistral AI, 2024), GPT-3.5 (gpt-3.5-turbo) (Brown et al., 2020), and GPT-4 (gpt-4-0613) (OpenAI, 2023).
@@ 4.3 Results
Detailed feedback from an external theorem prover effectively guides LLMs in verifying and refining explanations for NLI.
@@
Explanation length/complexity impacts formalisation and verification.
The e-SNLI dataset, which includes only a single explanatory sentence per example, shows the best overall performance. In contrast, the multiple-choice question answering datasets, QASC and WorldTree, exhibit comparatively lower performance. QASC typically contains 2 explanatory sentences, while WorldTree ranges from 1 to 16 sentences. As the number of explanatory sentences increases, so does the complexity of the logical reasoning required. Models show lower refinement performance in WorldTree when compared to e-SNLI and QASC, with only 3%, 5%, and 5% of Llama-70b, Mixtral-8x7b, and Mistral-small explanations being refined in WorldTree. Meanwhile, 29% and 35% of explanations are refined by GPT-3.5 and GPT-4 in WorldTree, respectively. This process involves synthesising multiple explanatory sentences to fulfill sub-goals, which must then be integrated to meet the overall hypothesis goal.
Iterative and categorical refinement can monotonically reduce syntactic errors in autoformalisation.
@@
Figure 8: Human evaluation of refined explanations in terms of factuality and triviality.
@@ 4.4 Ablation Study
We conducted an ablation study to further evaluate and disentangle the impact of autoformalisation on performance. To this end, we adopted GPT-4 exclusively for the autoformalisation component, while retaining the original models for explanation refinement and proof strategy generation. As shown in Figure 6, integrating GPT-4 for autoformalisation led to a significant increase in the number of explanations successfully refined across all models. For instance, Llama2-70b with GPT-4 as the formalisation component refined explanations from 7% to 65% in the e-SNLI dataset. For the multiple-choice question answering dataset, GPT-3.5 showed a relatively smaller increase from 44% to 48% and from 29% to 34%. Despite these improvements, a performance gap persists between GPT-4 and the other models, which is attributed to GPT-4’s superior symbolic reasoning capabilities required for explanation refinement from the identified logical errors.
Explanations are progressively made more complete and consistent through iterative refinement.
In order to deliver step-wise logical consistency, explanations need to be made complete and self-contained, leading to the introduction of additional explanatory sentences, which increases the total number of suggested proof steps. Therefore, we further evaluated how the proof steps vary when the total number of suggested proof steps increases, contrasting both refined and unrefined cases. Figure 7 illustrates this trend. In general, all models show a positive trend, as the total suggested proof steps increase, the average number of proof steps processed by the proof assistant also increases. Models like Mistral-small and GPT-3.5 tend to suggest more proof steps to accomplish the logical goal, which can result in some redundant steps, such as the significant pulse shown in Figure 7(c). For unrefined explanations, as shown in Figure 7(d), 7(e) and 7(f), the progression is steadier but retains a positive trend, where the models generally suggest more proof steps in response to the additional explanatory sentences introduced to correct a logical error identified from the erroneous step. We analysed the correlation between average successful explanatory sentences and total planned sentences in proofs, detailed in Appendix A.3. Examples of refined and unrefined explanations are in Appendix A.5.
@@ 4.5 Factual Errors and Trivial Explanations
In addition to evaluating the logical validity of explanations, we also conducted a human evaluation of the refined explanations, considering factual correctness and explanation triviality for the two best-performing models (GPT-3.5 and GPT-4). This evaluation focused on two questions: “Are the refined explanatory sentences factually correct?” and “Is the explanation trivial, merely repeating or paraphrasing the content of the premise and hypothesis to achieve logical validity?”. As illustrated in Figure 8, our findings indicate that all refined explanations in the e-SNLI and WorldTree datasets are consistent with commonsense knowledge. In the QASC dataset, 2.27% and 1.82% of the explanation refined by GPT-3.5 and GPT-4 contain sentences misaligned with true world knowledge. We found that the majority of these errors result from over-generalisation, such as the sentence All tetrapods are defined to have four limbs, which inaccurately includes snakes.
@@ 5.1 LLMs Self-Refinement from External Feedback
Self-refinement of LLMs has demonstrated promising effectiveness in generating faithful and trustworthy responses (Pan et al., 2023b). The use of external feedback to guide LLMs has been extensively studied (Yu et al., 2023; Akyurek et al., 2023; Olausson et al., 2024a). Previous work such as Peng et al. (2023) have employed facts retrieved from external knowledge bases as sources of feedback, while Paul et al. (2024) developed a critic model to provide feedback for reasoning refinement. Additionally, Nathani et al. (2023) have explored the use of feedback models for automated feedback generation. Various works have also investigated tasks related to code generation (Chen et al., 2023; Olausson et al., 2024b) and the creation of either synthetic or expert-written logical natural language expressions (Olausson et al., 2023). Quan et al. (2024) use a differentiable logic reasoner for verifying and refining explanations via abductive reasoning, improving logical consistency in ethical NLI tasks. This paper focuses on the automated verification and refinement of natural language explanations created by human annotators in NLI tasks. Our method leverages feedback from external solvers to iteratively refine explanations, which require specific modelling interventions such as extracting the exact erroneous steps from the theorem prover to effectively refine logical errors in the explanatory sentences.
5.2 Explanation Generation
Existing work has explored robust and effective approaches for multi-hop reasoning tasks in explanation generation (Thayaparan et al., 2021; Valentino et al., 2022b; Neves Ribeiro et al., 2022). In prior research, metrics such as Mean Average Precision (MAP) (Valentino et al., 2022a) have been employed to assess the ranking of facts in explanation generation tasks against gold-standard explanations. Although these metrics effectively measure precision relative to these standards, they inadequately capture the logical consistency and completeness of the explanations generated. Such shortcomings are particularly critical in tasks that require not only factual accuracy but also coherence and inferential soundness, as in natural language inference and explanation generation. Our proposed metrics address this gap by incorporating assessments of logical validity. Although some metrics have been proposed to manually evaluate the logical validity of explanations (Valentino et al., 2021; Yuan et al., 2024), such as non-redundancy or logical errors, these require significant effort from domain experts in formal languages. In this work, we use human-annotated explanations as a foundational dataset to detect and correct logical discrepancies, offering a framework adaptable for automatically enhancing both the precision and logical integrity of outputs across multi-step inference tasks.
@@ 5.3 Autoformalisation
Autoformalisation refers to the process of translating natural language descriptions into symbolic representations. Research in this area has included the formalisation of mathematical proofs (Cunningham et al., 2022; Wu et al., 2022; First et al., 2023; Jiang et al., 2023), and efforts to transform natural language sentences into logical forms using LLMs (Pan et al., 2023a; Olausson et al., 2023; Jiang et al., 2024b; Dalal et al., 2024). However, contextual information is frequently lost when sentences are translated in these logical frameworks. To mitigate semantic loss during the transformation process, we leverage Neo-Davidsonian event semantics, which aims to maximise the preservation sentence-level content. This representation paradigm can facilitate a more systematic content-preserving translation to logical forms, which is more independent from particular choices of representation schema.
6 Conclusion
In this work, we present a novel neuro-symbolic framework, Explanation-Refiner, which integrates LLMs and theorem provers for automatic verification and refinement of natural language explanations through iterative cycles. Extensive experiments on textual entailment and multiple-choice QA tasks showed improved logical validity of human-annotated explanations. We investigated the model’s performance from simple to complex explanatory/sentence structures and introduced a method to prevent the loss of semantic information in autoformalisation tasks with error correction. In future work, we aspire to enhance the framework’s robustness towards complex and unstructured explanations with fewer iterations required to improve the model’s efficiency.
Limitations
While this work have demonstrated significant improvements in terms of enhancing the logical consistency of explanations, the connection between logical consistency and AI safety still needs further investigation. While the idea of using formal solvers in conjunction with LLMs delivers a promise avenue to improve the consistency of reasoning within LLMs, these methodologies need to be further developed and critically assessed as a mechanism which can provide guarantees of correctness, consistency and completeness within critical application domains.
@@ A.4 Prompts
Temperature settings were adjusted to 0 for GPT-3.5 and GPT-4, and to 0.01 for Llama2-70b, Mixtral-8x7b, and Mistral-small, aiming to achieve both determinism in the output and effective code generation for theorem prover.
@@ A.5 Examples of Explanation Refinement
Table 1 shows an example from the e-SNLI dataset of how the explanation changes after each iteration. Figures 21, 22, and 23 illustrate the Isabelle/HOL theory code changes during the refinement process. Table 2 with Figures 24, 25, and 26 also show another example of how the explanation is refined after each iteration.
Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step. More examples can be found at https://github.com/neuro-symbolic-ai/explanation_refinement.
A.6 Datasets and Theorem Prover
The datasets used in our experiments, including samples from e-SNLI (Camburu et al., 2018), QASC (Khot et al., 2019), and WorldTree (Jansen et al., 2018), are all sourced from open academic works. We employed Isabelle as the theorem prover, which is distributed under the revised BSD license. Additionally, the TCP client used for the Isabelle server (Shminke, 2022) is licensed under Apache-2.0.